OpenStudy (anonymous):

Find the mass? Find the mass of acetic anhydride given: Density=1.08 g mL-1 Volume= 4mL (Molar mass of acetic anhydride=102.09 g mol-1 , but not sure if that is needed)

OpenStudy (nincompoop):

what is the formula for DENSITY?

OpenStudy (anonymous):

I realize that it's D=MV and my answer is 4.32, but I was just confused with the g/mol and the ml, if there is another step I needed

OpenStudy (nincompoop):

\[\rho = \frac{ m }{ v }\]

OpenStudy (anonymous):

I'm looking for mass

OpenStudy (anonymous):

not density

OpenStudy (abmon98):

1.08*4=4.32

OpenStudy (nincompoop):

\[m = \rho * v = 1.08 \frac{ g }{ mL } \times 4 mL\]

OpenStudy (nincompoop):

mL cancels out, and you are left with g as unit significant figure = 1

OpenStudy (anonymous):

You know, that would've been clear if I could read my own hand writing, thank you!